OpenStudy (anonymous):

Evaluate 7^log(4).... please help! (:

OpenStudy (anonymous):

that's the common log right?

OpenStudy (anonymous):

Correct (:

OpenStudy (anonymous):

sorry... something wrong

OpenStudy (anonymous):

Okay I saw what you did.... thank you so much <3

OpenStudy (anonymous):

but that was wrong....

OpenStudy (anonymous):

Did the log need to go in front like this.... log(4)=7?

OpenStudy (anonymous):

i don't think so. but my work was SweetlyFlawed and that's why I removed it... :)

OpenStudy (anonymous):

Ah, I see, terribly sorry to have misunderstood

OpenStudy (anonymous):

sorry :(

OpenStudy (anonymous):

Quite alright love, it is a mean problem indeed. (;

OpenStudy (anonymous):

If you can give the final answer, then maybe we can help you. There are many options possible, but the log ain't going to get removed

OpenStudy (anonymous):

I don't understand....

OpenStudy (anonymous):

Answer's possible a)7^log(4) = 7^ (2log2) b)7^log(4) = 4^(log 7)= 2^(2log 7) The log in the final answer remains. So I asked if you know the final answer of this question so that we can help you to get that answer

OpenStudy (anonymous):

But I do not know the final answer and I'm having trouble setting up the problem to solve it.

OpenStudy (anonymous):

you just want a number out of this?

OpenStudy (anonymous):

if so, here you are. 7^log(4) = 3.2270056

OpenStudy (anonymous):

\[7^{\log(4)}\] like that? calculator for this one

OpenStudy (anonymous):

No I didn't just want an answer... I wanted some help to figure it out adn explain the steps they took to recieve that answer...

OpenStudy (anonymous):

what was the answer?